# Think Together and Work Better: Combining Humans’ and LLMs’ Think-Aloud Outcomes for Effective Text Evaluation

### v1

@@ 5.3.2. Humans have strengths in internal quality
When comparing InteractEval’s performance using checklists based on LLM-only TA (SL-TA and ML-TA) with human-only TA (SH-TA and MH-TA) in both single and multiple settings according to Figure 5, human-only TA outperforms LLM-only TA in internal quality (Coherence and Fluency) of the generated texts. This suggests that humans excel in identifying attributes related to structure, semantic connections, and readability, which are critical to internal text quality. Their flexible thinking skills (Blattberg and Hoch, 1990) and innate understanding of universal grammar (Chomsky et al., 2023; Montague et al., 1970) allow them to generate higher-quality attributes that account for the hierarchical structure of texts (Pantcheva, 2023), a task that is more challenging for LLMs.
Qualitative analysis supports this conclusion. In the analysis, (A) refers to the attributes generated by humans and (B) refers to those generated by LLMs. For Coherence, humans created more quantitative and explicit attributes, making an evaluator easier to measure consistently. E3 commented, ”(A) offers simple, straightforward evaluations with easily quantifiable items, while (B) includes more qualitative, subjective items.” Additionally, humans focused more on logical order and organization. E1 noted, ”Attribute (A) emphasizes clear flow, meaningful connections, and sentence links for efficient information delivery, while (B) focuses more on comprehension but is less efficient.” For Fluency, humans provided attributes in more detail, while LLMs took a broader approach. LLM-generated attributes sometimes confused experts due to overlap with other dimensions. E1 remarked, ”Attribute (A) offers detailed and specific requirements, like including at least two interrogatives, while (B) provides more general guidelines.” E2 noted, ”Section (B) had many ambiguous items and concepts related to consistency.”
Additionally, the attribute similarities shown in Figure 7 support these findings. Human-generated attributes for internal quality are less lexically and semantically similar compared to LLM-generated ones, suggesting that humans’ TA produces more diverse attributes, allowing checklists to assess summaries across a wider range of features.
@@
In terms of external alignment (Consistency and Relevance), our framework using checklists based on LLM-only TA (SL-TA and ML-TA) shows superior performance-or minimal differences—compared to those based on human-only TA (SH-TA and MH-TA). This suggests that LLMs excel in identifying attributes related to external alignment, where both the source text and the summary must be considered simultaneously. The task requires processing large amounts of information, where LLMs’ vast context window and expansive memory excel (Pantcheva, 2023; Achiam et al., 2023; Wu et al., 2023). LLMs provide consistent attributes even when handling large data sets, adhering to their pre-trained policies (Blattberg and Hoch, 1990; Pantcheva, 2023).
Qualitative analysis supports these findings. For Consistency, LLMs explain the attributes in more detail and cover a broader scope than humans. E1 noted, ”Attribute (B) describes each point with more detail,” and E4 added, ”Attribute (B) fully explains more details compared to (A).” Additionally, LLMs provide systematic guidelines for detecting hallucinations, while human experts only provide attributes to check for false content. E2 commented, ”(B) gives detailed guidelines for identifying hallucinations, while (A) only shows how to detect them.” For Relevance, humans tend to create confusing attributes that could apply to multiple dimensions. One expert mentioned being confused, thinking that certain attributes were Coherence: ”In (A), some sentences could be interpreted as a coherence issue.” E4 also noted, ”Sometimes (A) felt like it belonged in another category.” LLMs, on the other hand, clearly reflect the rubric’s definition of Relevance and generate attributes covering a broader scope. Three out of four experts agreed on this. E3 said, ”(B) covers more areas and provides more accurate and detailed descriptions for each item according to the rubric.”
As a result, LLM-generated attributes lead to better checklist construction and improved evaluation performance in the dimensions related to external alignment, as shown in Figure 5. Moreover, Figure 7 shows that LLM-generated attributes for external alignment are less lexically and semantically similar compared to those created by humans, indicating that LLM-TA produces more diverse attributes, enabling checklists to assess summaries across a wider range of features.
@@ 6. Discussion
Figure 8. Comparison of evaluation performance averaged across four dimensions using two correlation metrics (Spearman and Kendall-tau). Single refers to the average performance of InteractEval with checklists based on SL-TA and SH-TA, Multiple refers to the average performance with checklists based on ML-TA and MH-TA, and Combination represents the performance using checklists based on Comb-TA.
@@ 6.1. Major Findings
The goal of this study is to develop a text evaluation framework, InteractEval, using checklists built from attributes generated through the combined TA processes of humans and LLMs. Our key findings are as follows. First, InteractEval outperforms both traditional non-LLM and LLM-based baselines in terms of correlation with human-assigned scores. Analysis of attribute topics and similarities shows that this improvement stems from reducing biases and enhancing fine-grainedness by incorporating both human and LLM perspectives, as well as rubrics and sample texts, supporting RQ 1.
Secondly, the TA process encourages divergent thinking in LLMs, generating a wider range of text attributes, which form the basis of the checklists. This is confirmed by the strong performance of checklists based on the attributes from single-LLM TA, supporting RQ 2.
Finally, in response to RQ 3, combining the TA outcomes of humans and LLMs mitigates biased attributes and increases their fine-grainedness, leading to more refined checklists and improved evaluation performance. Additionally, humans and LLMs exhibit distinct strengths in text evaluation. Humans excel in assessing internal text quality, where flexible thinking and an innate understanding of grammar are crucial. In contrast, LLMs are more effective at detecting external alignment, thanks to their ability to process large amounts of information consistently.
Therefore, combining the complementary outcomes of both humans’ and LLMs’ TA in text evaluation enhances overall performance, as shown in Figure 6(b), compared to using attributes from either humans or LLMs alone.
@@ 6.2. Implications
This study offers practical and research-related implications. As previously noted, text generation is rapidly expanding alongside advancements in LLMs. At the same time, the need for accurate methods to evaluate the quality of generated texts is gaining significant interest. InteractEval provides a cost-effective metric that shows reliable correlations with human assessments. Similarly, in research areas where reliable metrics for evaluating LLMs’ text generation abilities are scarce, our framework can serve as a useful testing method. Furthermore, in the context of rapidly increasing machine-centered automation, including LLMs, this study underscores the importance of human high-level thinking abilities and suggests a direction for effective human-machine interaction.
7. Limitations
@@
In this study, we identified three main limitations. First, we tested our framework using the SummEval dataset (Fabbri et al., 2021) because it includes the two key features we focus on: internal quality and external alignment. As a result, our findings may not generalize to other datasets or text generation tasks, requiring further research. Secondly, during the experiment, we observed differences in the capabilities of humans and LLMs when constructing attributes for checklists. However, because comparing these capabilities was not the main focus of this study, we did not explore them in depth. Lastly, to ensure a fair comparison with baselines that primarily use GPT models, we did not use smaller LLMs, such as Llama-3.1-8B-Instruct, as evaluators. Future studies could include smaller LLMs to assess the proposed methods’ capabilities.
8. Conclusion and Future Work
Our work distinguishes itself from other LLM-based text evaluation studies. We introduce InteractEval, a framework that integrates human expertise and LLMs using the think-aloud (TA) method to generate attributes for checklist-based text evaluation. Our findings demonstrate that combining the strengths of humans and LLMs improves evaluation performance across the four dimensions, Coherence, Fluency, Consistency, and Relevance, highlighting the value of their collaboration (RQ1).
The study also emphasizes the effectiveness of TA in promoting divergent thinking in both humans and LLMs, helping to identify a broader range of attributes for each evaluation dimension (RQ2). This approach results in more comprehensive and refined checklists, ultimately enhancing the quality of text evaluation.
Additionally, we show that leveraging both human experts and LLMs together produces the best outcomes, reinforcing the importance of their combination (RQ3). Humans excel in identifying attributes related to internal quality (Coherence and Fluency), while LLMs outperform in external alignment (Consistency and Relevance).
Future research should assess the robustness and generalizability of this checklist-based approach, which is built on attributes generated by humans and LLMs through the TA process, using a variety of real-world text generation datasets. Additionally, exploring the strengths and weaknesses of humans and LLMs across different tasks could provide further insights. Ultimately, as the title suggests, this study demonstrates that when LLMs and humans think together, they can work better in text evaluation.
Appendix A Example of System Messages and Predefined Template for LLMs TA
A.1. System Message
You are an assessor who rate {Dimension} of the summary of a news article. You must read the summary samples and rubric for rating {Dimension}. While reading, write down your thoughts on what to keep in mind when evaluating the {Dimension} of the text, referring to the rubric. Avoid mentioning content specific to the samples. Instead, focus on general considerations for evaluating the {Dimension} of any summary based on the provided rubric.
A.2. Predefined Template
Rubric of {Dimension}:
{Rubric}
¡Sample 1¿
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
¡End of Sample 1¿
¡Sample 4¿
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
¡End of Sample 4¿
Please provide each consideration in a JSON format with the key of consideration index and the value of the consideration. Do not name each consideration. For example, {1: ”consideration 1”, 2: ”consideration 2”, 3: ”consideration 3”, 4: ”consideration 4”, 5: ”consideration 5”, 6: ”consideration 6”,…}
Appendix B Example of System Messages and Predefined Template for Component Extraction
B.1. System Message
1. The components should be related to {Dimension} defined below.
2. Remove any components that are ambiguous or redundant.
3. The number of the components should not be more than 5.
Definition of {Dimension}: {Definition}
B.2. Predefined Template
Sentences: {Attributes}
For example, [”component 1”, ”component 2”, ”component 3”, …]
Appendix C Example of System Messages and Predefined Template for Attributes Clustering
C.1. System Message
You can remove attributes that are redundant or confusing.
C.2. Predefined Template
Components: {Components}
Attributes: {Attributes}
Please, provide your answer in a JSON format with the keys of the components and the values of the list of attributes.
For example, {”component 1”: [”attribute 1”, ”attribute 2”], ”component 2”: [”attribute 3”, ”attribute 4”, ”attribute 5”], ”component 3”: [”attribute 6”], …}
Appendix D Example of System Messages and Predefined Template for Question Generation
The prompt template is inspired by the prompts introduced in (Lee et al., 2024b).
D.1. System Message
You are a key question generator for evaluating {Dimension}.
2. Based on these common features, construct a set of Yes-No questions related to each key component. These questions should be designed to effectively evaluate the accuracy and relevance of a summary of a news article.
D.2. Predefined Template
The definition of {Dimension} and the set of components and their attributes corresponding to {Dimension} are provided below.
Each component and question must satisfy the following conditions:
1. Each component should have one corresponding question.
2. Each question must be answerable with ‘Yes’ or ‘No’.
3. Each question must incorporate concepts from the key component.
5. Formulate questions so that a ‘Yes’ answer is a positive response.
# Definition
{Rubric of a dimension}
# Components and attributes
{Components and attributes}
Appendix E Example of System Message and Predefined Template for Sub-Question Generation
The prompt template is inspired by the prompts introduced in (Lee et al., 2024b).
E.1. System Message
E.2. Predefined Template
In this task, you need to create a question to evaluate the {Dimension} of the summary of the original document. The definition of {Dimension} and the questions corresponding to the key component of {Dimension} are provided below. Use them to generate sub-questions for each key question.
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ‘Yes’ or ‘No’.
5. Formulate sub-questions so that a ‘Yes’ answer is a positive response.
6. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones. For example, ask whether a summary avoids negative elements rather than contains them.
# Definition
{Rubric of a dimension}
# Components and corresponding questions
{Components and questions}
Please provide your answer in a JSON format with the key of components and the value of the list of sub-questions.
For example, {”component 1”: [”sub question 1”, ”sub question 2”,…] , ”component 2”: [”sub question 1”, ”sub question 2”, ”sub question 3”, …], ”component 3”: [”sub question 1”, ”sub question 2”, …], ”component 4”: [”sub question 1”, ”sub question 2”, ”sub question 3”, ”sub question 4”, …],…}
Appendix F Example of System Message and Predefined Template for Question Validation
F.1. System Message
You are a sub-question evaluator.
F.2. Predefined Template
In this task, you need to examine each sub-question for evaluating the {Dimension} of the summary of the original document. The definition of {Dimension} is provided below.
# Definition
{Rubric of a dimension}
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ‘Yes’ or ‘No’.
2. Each question must contain concepts of the {Dimension}.
5. Formulate questions so that a ‘Yes’ answer is a positive answer.
7. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones. For example, ask whether a summary avoids negative elements rather than contains them.
# Sub-questions {Sub-questions}
For example, [”sub question 1”, ”sub question 2”, ”sub question 3”, …]
Appendix G Open-ended questions for qualitative analysis

### v2

@@ 5.3.2. Humans have strengths in internal quality
When comparing InteractEval’s performance using checklists based on LLM-only TA (SL-TA and ML-TA) with human-only TA (SH-TA and MH-TA) in both single and multiple settings, human-only TA outperforms LLM-only TA in internal quality (Coherence and Fluency) of the generated texts (see Figure 9), suggesting that humans excel in identifying attributes related to structure, semantic connections, and readability, which are critical to internal text quality. Their flexible thinking skills (Blattberg and Hoch, 1990) and innate understanding of universal grammar (Chomsky et al., 2023; Montague et al., 1970) allow them to generate higher-quality attributes that account for the hierarchical structure of texts (Pantcheva, 2023), a task that is more challenging for LLMs.
Qualitative analysis corroborates this finding. For ease of exposition, in this analysis, (A) refers to the attributes generated by humans and (B) refers to those generated by LLMs, during the TA process. For Coherence, humans created more quantitative and explicit attributes, making the qualitative analysis participants (E1-E4) easier to measure consistently. E3 commented, ”(A) offers simple, straightforward evaluations with easily quantifiable items while (B) includes more qualitative, subjective items.” Additionally, humans focused more on logical order and organization. E1 noted, ”(A) emphasizes clear flows, meaningful connections, and natural sentence links for efficient information delivery while (B) focuses more on comprehension but is less efficient.” For Fluency, humans provided attributes in more detail while LLMs took a broader approach. LLM-generated attributes sometimes confused qualitative study participants due to overlap with other dimensions. E1 remarked, ”(A) offers detailed and specific requirements, like including at least two interrogatives while (B) provides more general guidelines.” E2 noted, ”(B) had many ambiguous items and concepts related to consistency.” Additionally, the attribute similarities shown in Figure 11 support these findings. Human-generated attributes for internal quality are less lexically and semantically similar compared to LLM-generated ones, suggesting that humans’ TA produces more diverse attributes, enabling checklists to assess summaries across a wider range of features.
@@
In terms of external alignment (Consistency and Relevance), our framework using checklists based on LLM-only TA (SL-TA and ML-TA) shows superior performance or at least minimal differences, compared to those based on human-only TA (SH-TA and MH-TA), suggesting that LLMs excel in identifying attributes related to external alignment, where both the source text and the summary must be considered simultaneously. The task requires processing a large amount of information, where LLMs’ vast context window and expansive memory can excel (Pantcheva, 2023; Achiam et al., 2023; Wu et al., 2023). LLMs are capable of providing a reliable set of attributes even when handling large data, adhering to their pre-trained policies (Blattberg and Hoch, 1990; Pantcheva, 2023).
Qualitative analysis further supports these findings. For Consistency, LLMs explain the attributes in more detail and cover a broader scope than humans. E1 noted, ”(B) describes each point with more detail,” and E4 added, ”(B) fully explains more details compared to (A).” Additionally, LLMs provide systematic guidelines for detecting hallucinations while human experts only provide attributes to check for false content. E2 commented, ”(B) gives detailed guidelines for identifying hallucinations while (A) only shows how to detect them.” For Relevance, humans tend to create confusing attributes that could be mapped to multiple dimensions. One expert mentioned being confused, thinking that certain attributes were Coherence: ”In (A), some sentences could be interpreted as a coherence issue.” E4 also noted, ”Sometimes (A) felt like it belonged in another category.” LLMs, on the other hand, clearly reflected the rubric’s definition of Relevance and generated attributes covering a broader scope. Three out of four experts agreed on these points. E3 said, ”(B) covers more areas and provides more accurate and detailed descriptions for each item, on the basis of the rubric.”
As a result, LLM-generated attributes lead to better checklist construction and improved evaluation performance in the external alignment dimensions, as shown in Figure 9. Moreover, Figure 11 shows that LLM-generated attributes for external alignment are less lexically and semantically similar compared to those created by humans, indicating that LLM-TA produces more diverse attributes, enabling checklists to assess summaries across a wider range of features.
@@ 6.1. Major Findings
Figure 12. Comparison of evaluation performance averaged across four dimensions using two correlation metrics (Spearman’s Rho ($\rho$) and Kendall’s Tau ($\tau$)). Single refers to the average performance of InteractEval with checklists based on SL-TA and SH-TA, Multiple refers to the average performance with checklists based on ML-TA and MH-TA, and Combination represents the performance using checklists based on Comb-TA.
@@
In this study, we propose InteractEval, a text evaluation framework that uses checklists derived from attributes generated through Think Aloud (TA) process based on human-machine collaboration. It assesses the framework’s effectiveness compared to traditional and state-of-the-art baselines on a text summarization task, examines its transferability to a different task (essay scoring), and explores the potential synergistic effects of combining TA outcomes from humans and large language models (LLMs). Our key findings can be summarized as follows. First, InteractEval outperforms both traditional non-LLM and LLM-based baselines on evaluating summarized texts (Table 9) and essays (Figure 7) in relation to correlating with the human-assigned ground-truth scores provided by the dataset. Particularly, analysis of attribute topics and similarities on the summary evaluation task shows that this improvement stems from reducing biases and enhancing fine-grainedness by incorporating both human and LLM perspectives, as well as rubrics and sample texts, supporting RQ 1.
Secondly, the prompt engineering procedure employed for LLMs in this study, which is akin to the TA process employed for human experts, encourages LLMs to generate a wider range of text attributes, which form the basis of the checklists. This argument has been confirmed by the strong performance of checklists based on the attributes from single-LLM TA, supporting RQ 2. In particular, considering that TA with each LLM demonstrates high performance in different dimensions, we conclude that combining the TA results from multiple LLMs complements each LLM’s outcome and produces more reliable and refined checklists.
Finally, in response to RQ 3, combining the TA outcomes of humans and LLMs mitigates biased attributes and increases their fine-grainedness, leading to more refined checklists and improved evaluation performance. Specifically, humans and LLMs exhibit distinct strengths in text summary evaluation. Humans excel in assessing internal text quality, leveraging flexible thinking and an innate understanding of grammar. For example, attributes generated from human-only TA effectively capture the coherent and hierarchical structure of text (Pantcheva, 2023), closely aligning with internal quality dimensions such as Coherence and Fluency. Consequently, more refined checklists can be designed through explicit and detailed attributes, resulting in more reliable evaluation performance. These findings highlight that human’s empirical knowledge and practical skills surpass LLMs in capturing composition and organization features, such as semantic coherence and readability, consistent with a prior study (Dhillon et al., 2024). In contrast, LLMs excel at detecting external alignment due to their ability to process large volumes of information consistently. Attributes generated from LLM-only TA effectively capture the semantic and structural connections between summary and source texts, closely related to external alignment dimensions such as Consistency and Relevance. These attributes improve checklist design by covering broader scopes and providing systematic guidelines for detecting hallucinations, underscoring LLMs’ ability to efficiently manage large amounts of textual information and consistently compare two texts, which are in support of prior studies (Blattberg and Hoch, 1990; Wang et al., 2024a). Therefore, combining the outcomes of both humans’ and LLMs’ TA in checklist design leverages their complementary strengths, resulting in an optimal checklist and improved text evaluation performance for tasks such as text summarization (Figure 12), compared to using attributes from humans or LLMs alone. Additionally, the enhanced performance of InteractEval on the essay evaluation dataset underscores the transferability of its combinatory effect. This finding further highlights the advantages of human-AI collaboration, building on prior studies across various domains (Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023; Revilla et al., 2023; Cohen et al., 2023).
@@ 6.2. Implications
This study offers practical as well as research implications. As previously noted, text generation is rapidly expanding alongside advancements in LLMs. At the same time, the need for accurate methods to evaluate the quality of generated texts is gaining significant interest across various industry domains including education. InteractEval is a framework designed to answer the call. The framework can be utilized in various industry settings where the quality assessment of generated texts is critical. Similarly, for those research areas exploring the use of LLMs for text generation but lacking reliable metrics to evaluate the quality of the generated texts, our framework can serve as a useful testing tool. Moreover, our study empirically validates the impact of human-LLM collaboration during the think-aloud process, which constitutes the first stage of checklist design. By doing so, it suggests both the necessity and potential for future research to explore how collaboration between humans and LLMs in subsequent stages could further amplify this effect. Beyond the evaluation performance results, InteractEval demonstrates how humans and LLMs can work together to more reliably evaluate summaries and essays, anticipating that their synergistic potential will grow alongside advancements in LLM technology.
7. Limitation and Future Work
@@
For this study, we have identified three main limitations. First, we tested our framework on a text summary evaluation task (SummEval (Fabbri et al., 2021)) and an essay evaluation task (ELLIPSE (Crossley et al., 2023)). However, our findings may not generalize to other types of text-related tasks, such as question-and-answering, image description, or text generation, as those tasks have different requirements for evaluation. Thus, further research is needed to explore the applicability of human-AI collaboration to text-related tasks beyond the evaluation of summaries and essays. Secondly, to ensure a fair comparison with the baselines that primarily use GPT models, we did not use smaller LLMs, such as Llama-3.1-8B-Instruct, as evaluators. Future studies could include the small-sized LLMs to assess the proposed methods’ capabilities. Lastly, the two datasets used in this study are based in English, leaving room to examine the applicability of our approach to text evaluation tasks in other language settings.
8. Conclusion
Our work distinguishes itself from other LLM-based text evaluation studies by paying attention to the potential synergetic effects that might be achieved by combining the strengths of humans and LLMs. To explore the understudied potential of human-AI collaboration in the context of text evaluation, we introduce InteractEval, a framework that integrates human expertise and LLMs using the concurrent think-aloud (TA) method to generate attributes for checklist-based text evaluation. Our findings demonstrate that combining the strengths of humans and LLMs significantly enhances text evaluation performance, highlighting the synergistic effect of their collaboration. Our study also emphasizes the effectiveness of TA in promoting divergent thinking in humans and divergent generation in LLMs, helping to identify a broader range of attributes, providing a clear explanation of the mechanism responsible for the superior performance outcomes observed. Furthermore, we empirically demonstrate that human experts and LLMs have their distinct but combinatory strengths. Humans excel in identifying attributes related to internal quality (Coherence and Fluency) whereas LLMs outperform in external alignment (Consistency and Relevance). Ultimately, as the title suggests, this study demonstrates that when humans and and LLMs collaborate, they can work better in text evaluation, opening up new horizons for effective human-machine collaboration.
Acknowledgements.
This research was supported by Chunjae Education Inc. (Project Number G01250064).
Appendix
Appendix A Checklists of three rating dimensions in SummEval dataset (Table 12, 13, and 14)
Appendix B Checklists of seven rating dimensions in ELLIPSE dataset (Table 15, 16, 17, 18, 19, 20, and 21)
Appendix C Information of human experts who participated in the TA process for ELLIPSE dataset (Table 22)
Appendix D Comparison of distributions in ELLIPSE dataset (Figure 13).
Figure 13 shows the score distributions for the entire ELLIPSE dataset, the first sampled dataset, and second sampled dataset across all dimensions. As shown, the sampled datasets preserve distributions nearly identical to those of the original dataset.
Appendix E Evaluation performance on ELLIPSE dataset (Table 23 and 24).
Table 23 and 24 present the text evaluation performance on the ELLIPSE dataset. The table demonstrates that InteractEval’s performance in both the first and second trials, using GPT-4, surpasses the baselines across all dimensions in terms of Spearman’s Rho and Kendall’s Tau correlations, except for the Spearman’s Rho correlation in Cohesion.
Appendix F Open-ended questions for qualitative analysis